# A New Perspective on the Diffuse Gamma-Ray Emission Excess

### v1

@@ I Introduction
Recently, the Large High-Altitude Air Shower Observatory (LHAASO) published the gamma-ray source catalog of the TeV-PeV energy band Cao et al. (2024a) and the measurement of DGE in the 10-1000 TeV energy range by masking the source region Cao et al. (2023). The measured DGE flux in the inner Galaxy region is about 3 times higher than the expectation from the CR-ISM interactions and about 2 times higher than that expected in the outer Galaxy region Zhang et al. (2023). On the other hand, neutrinos are produced alongside gamma rays during the CR-ISM interactions. The IceCube neutrino telescope has measured the high-energy neutrino flux from the Galactic plane in the energy range of 1-100 TeV Abbasi et al. (2023). Subtracting possible contribution of point sources from the total neutrino measurement and applying the masking method of LHAASO, Ref. Yan et al. (2024) argued that the gamma-ray flux associated with the neutrino flux is consistent with that predicted by the CR-ISM interactions. This result also supports that the TeV DGE excess is mainly contributed by leptonic processes [^1] Due to the model-dependent nature of the neutrino flux measurement and the associated uncertainties, we cannot rule out the possibility that the DGE excess includes a minor hadronic component..
@@
This work aims to evaluate the contribution of LHAASO PWNe/halos leakage signals to the TeV DGE excess based on the two-zone diffusion model of electrons, which is a completely different perspective compared to attributing the DGE excess to unresolved extended gamma-ray sources Linden and Buckman (2018); Yan et al. (2024). Additionally, we give an estimate of the signal leakage from the Cygnus bubble. In Section II, we introduce the two-zone diffusion model, as well as the measurements of LHAASO PWNe/halos in the first LHAASO catalog and the related pulsar characteristics. We adopt the two-zone diffusion model to fit the LHAASO measurements and then predict the signal leakage. Section III presents our results, including the Galactic longitude profile and gamma-ray energy spectrum contributed by the leakage signals. Section IV provides some extended discussions of the results, followed by a summary and outlook in Section V.
@@ II.1 Two-zone model
We assume the electron propagation for the LHAASO PWNe/halos can be described by the diffusion-cooling equation:
@@
where $N$ is the electron number density and $E_{e}$ is the electron energy. The source term $Q$ is assumed to have a point-like characteristic, and the electron injection rate is assumed to follow the time profile of the pulsar spin-down luminosity as as $\propto\left(1+t/\tau\right)^{-2}$, where the spin-down time scale is set to be $\tau=10~{}\rm{kyr}$. The electron injection spectrum for each source is described by a power-law form. Hence, the source injection function is expressed as
@@ II.3 Fitting method
Figure 1: The surface brightness profile of LHAASO J1825-1337u with $E_{\gamma}=32$ TeV. The blue line represents the prediction of a Gaussian profile, with corresponding shaded area indicating the error band, assuming the parameter errors are independent of each other. The red line represents the expectation from the two-zone diffusion model. The gray shaded area indicates the fitting range.
@@ III Results
Figure 2: The gamma-ray intensity map of Galactic plane contributed by LHAASO PWNe/halos and Cygnus bubble in energy range 10-63 TeV. The signals from LHAASO PWNe/halos are modeled using a two-zone diffusion model. (a) The inner Galactic plane of $15^{\circ}<l<125^{\circ}$. (b) The outer Galactic plane of $125^{\circ}<l<235^{\circ}$. (c) The inner Galactic plane after masking the same region as the LHAASO DGE measurement. (d) The outer Galactic plane after masking the same region as the LHAASO DGE measurement.
@@
Figure 3: The comparison of Galactic profiles of the diffuse emission and model. The top panel is for the energy band 10-63 TeV, and the bottom panel is for 63-1000 TeV. The black points represent the measurements from LHAASO. The red line represents the prediction of the CR-ISM interaction modelYan et al. (2024). The blue line represents the sum of contributions from the CR-ISM interaction and the leakage signals of the LHAASO PWNe/halos and the Cygnus bubble. For sources with significant signal leakage, we have specifically marked their contributions to the DGE.
@@
Figure 4: Comparison of the measured DGE spectrum and model interpretation. The left panel shows the inner Galaxy region ($15^{\circ}<l<125^{\circ}$) and the right panel shows the outer Galaxy region ($125^{\circ}<l<235^{\circ}$). For each panel, the black points represent the LHAASO measurement. The red line represents the prediction of the CR-ISM interaction modelYan et al. (2024), the pink line represents the contribution from the leakage fluxes of LHAASO PWNe/Halo, and the black line represents the leakage flux of Cygnus bubble. The blue line is the sum of all these components.
@@ IV Discussion
We have demonstrated that the DGE excess in most regions can be interpreted by the signal leakage model under certain parameters. However, the gamma-ray flux predicted by the model is insufficient to match the data in the range of $l\approx 115^{\circ}-150^{\circ}$, even when considering model uncertainties, as shown in Figure 5. There are relatively few LHAASO PWNe/halos in this region, and the fluxes of LHAASO J0249$+$6022 and LHAASO J0359$+$5406 located in this region are relatively low. The excess gamma-ray emission in this region may be caused by some unresolved leptonic sources, such as indistinguishable halos. Therefore, the signal leakage from known sources and the contribution from unresolved sources should be considered complementary in explaining the DGE excess.
@@
Figure 5: Comparison of the Galactic longitude profiles of DGE by assuming different two-zone diffusion model parameters. The lines represent the total contribution from the CR-ISM interaction, the LHAASO PWNe/Halo, and the Cygnus bubble. The units of $D_{1}$ are cm2 s-1.

### v2

@@ I Introduction
Recently, the Large High-Altitude Air Shower Observatory (LHAASO) published the gamma-ray source catalog of the TeV-PeV energy band Cao et al. (2024a) and the measurement of DGE in the 10-1000 TeV energy range by masking the source regions Cao et al. (2023). The measured DGE flux in the inner Galaxy region is about 3 times higher than the expectation from the CR-ISM interactions and about 2 times higher than that expected in the outer Galaxy region Zhang et al. (2023). On the other hand, neutrinos are produced alongside gamma rays during the CR-ISM interactions. The IceCube neutrino telescope has measured the high-energy neutrino flux from the Galactic plane in the energy range of 1-100 TeV Abbasi et al. (2023). Subtracting possible contribution of point sources from the total neutrino measurement and applying the masking method of LHAASO, Ref. Yan et al. (2024) argued that the gamma-ray flux associated with the neutrino flux is consistent with that predicted by the CR-ISM interactions. This result supports that the TeV DGE excess is mainly contributed by leptonic processes [^1] Due to the model-dependent nature of the neutrino flux measurement and the associated uncertainties, we cannot rule out the possibility that the DGE excess includes a minor hadronic component..
@@
In this work, we introduce a novel approach by evaluating the contribution of leakage signals from LHAASO PWNe/halos to the TeV DGE excess based on the two-zone diffusion model of electrons. This perspective is entirely distinct from the previous attribution of the DGE excess to unresolved extended gamma-ray sources Linden and Buckman (2018); Dekker et al. (2024); Yan et al. (2024). Additionally, we give an estimate of the signal leakage from the Cygnus bubble. In Section II, we introduce the two-zone diffusion model, as well as the measurements of LHAASO PWNe/halos in the first LHAASO catalog and the related pulsar characteristics. We adopt the two-zone diffusion model to fit the LHAASO measurements and then predict the signal leakage. Section III presents our results, including the Galactic longitude profile and gamma-ray energy spectrum contributed by the leakage signals. Section IV provides some extended discussions of the results, followed by a summary and outlook in Section V.
@@ II.1 Two-zone model
We assume that the electron propagation for the LHAASO PWNe/halos can be described by the diffusion-loss equation:
@@
where $N$ is the differential electron number density at electron energy $E_{e}$, position $\bm{r}$ and time $t$. $D$ is the diffusion coefficient, $b\equiv|dE_{e}/dt|$ is the energy-loss rate due to electromagnetic radiation, and $Q$ is the source term.
The source is assumed to be point-like, and the electron injection rate is assumed to follow the time profile of the pulsar spin-down luminosity as as $\propto\left(1+t/\tau\right)^{-2}$, where the spin-down time scale is set to be $\tau=10~{}\rm{kyr}$. The electron injection spectrum for each source is described by a power-law form. Hence, the source injection function is expressed as
@@ II.3 Fitting method
$\chi^{2}=\sum_{i}^{n}\left(\frac{S(\theta,E_{\gamma})-S_{G}(\theta,E_{\gamma})}{\sigma_{G}^{i}}\right)^{2}$ (4)
where $i$ is the data sampling point, $n$ is the total number of sampling points, the angular distance between sampling points is $0.1^{\circ}$. $S_{G}(\theta,E_{\gamma})$ is surface brightness predicted by the Gaussian profile, $S(\theta,E_{\gamma})$ is surface brightness predicted by the two-zone diffusion model, and $\sigma_{G}^{i}$ is the point error based on Gaussian profile error, which is adjusted according to the sample size $n$. The morphology derived from the two-zone diffusion model is also convolved with the same PSF. The free parameters in the fitting process are $q_{0}$ and $\alpha$ defined in equation (2), which determine the normalization and spectral shape of the gamma-ray emission, respectively.
@@
Figure 1: The surface brightness profile of LHAASO J1825-1337u with $E_{\gamma}=32$ TeV. The blue line represents the prediction of the Gaussian profile, with corresponding shaded area indicating the error band, assuming the parameter errors are independent of each other. The red line represents the expectation from the two-zone diffusion model. The gray shaded area indicates the fitting range.
@@ III Results
Figure 2: The gamma-ray intensity map of Galactic plane contributed by LHAASO PWNe/halos and Cygnus bubble in energy range 10-63 TeV. The signals from LHAASO PWNe/halos are modeled using a two-zone diffusion model. (a) The inner Galactic plane of $15^{\circ}<l<125^{\circ}$. (b) The outer Galactic plane of $125^{\circ}<l<235^{\circ}$. (c) The inner Galactic plane after masking the same region as the LHAASO DGE measurement. (d) The outer Galactic plane after masking the same region as the LHAASO DGE measurement. The white regions in (c) and (d) represent the masked regions.
@@
Figure 3: The comparison of Galactic profiles of the diffuse emission and the model. (a) In the energy range of $10-63$ TeV. (b) In the energy range of $63-1000$ TeV. The black points represent the measurements from LHAASO. The red line represents the prediction of the CR-ISM interaction model Yan et al. (2024). The blue line represents the sum of contributions from the CR-ISM interaction and the leakage signals of the LHAASO PWNe/halos and the Cygnus bubble. For sources with significant signal leakage, we have specifically marked their contributions to the DGE.
@@
Figure 4: Comparison of the measured DGE spectrum and model interpretation. (a) In the inner Galactic plane of $15^{\circ}<l<125^{\circ}$. (b) In the the outer Galactic plane of $125^{\circ}<l<235^{\circ}$. For each panel, the black points represent the LHAASO measurement. The solid red line represents the prediction of the CR-ISM interaction model Yan et al. (2024), the dotted pink line represents the contribution from the leakage fluxes of LHAASO PWNe/Halo, and the black dashed line represents the leakage flux of Cygnus bubble. The blue line is the sum of all these components.
@@ IV Discussion
The uncertainties of the model arise from the parameters of the electron injection spectrum and the parameters of electron propagation. The injection parameters ($q_{0}$ and $\alpha$) are constrained by the LHAASO spectral measurements of the sources. The propagation parameters are mainly constrained by the morphology measurements of the sources; however, LHAASO currently provides only measurements of Gaussian extents for most sources, resulting in weaker constraints on the propagation parameters. Therefore, the uncertainties of the model primarily stem from the propagation parameters, especially $D_{1}$ and $r_{*}$, which determine the amount of signal leakage.
@@
We have demonstrated that the DGE excess in most regions can be interpreted by the signal leakage model under certain parameters, e.g., $D_{1}\approx 5\times 10^{27}$ cm2 s-1 and $r_{\star}\approx 20-25$ pc. However, the gamma-ray flux predicted by the model is insufficient to match the data in the range of $l\approx 115^{\circ}-150^{\circ}$, even when considering model uncertainties, as shown in Figure 5. There are relatively few LHAASO PWNe/halos in this region, and the fluxes of LHAASO J0249$+$6022 and LHAASO J0359$+$5406 located in this region are relatively low. The excess gamma-ray emission in this region may be caused by some unresolved leptonic sources, such as indistinguishable halos. Therefore, the signal leakage from known sources and the contribution from unresolved sources should be considered complementary in explaining the DGE excess.
@@
We note that some studies on the joint analysis of the LHAASO diffuse gamma-ray measurement and IceCube neutrino measurement suggest a hadronic origin for DGE Fang and Murase (2023); Shao et al. (2023). It is crucial to emphasize that a consistent treatment of both datasets is essential for a robust joint analysis Yan et al. (2024). The neutrino flux measured by IceCube includes contributions from both the CR sea and hadronic point sources. To correlate the neutrino flux measured by IceCube with the diffuse gamma-ray data from LHAASO, the contributions from hadronic sources should be subtracted. Furthermore, the neutrino flux should be derived from the same region of interest as the LHAASO DGE analysis. If the same masking method used by LHAASO is not performed to the IceCube data, or if the contributions from hadronic point sources are not subtracted, it could lead to an overestimation of the neutrino flux, potentially resulting in the conclusion that the DGE excess has a hadronic origin.
Figure 5: Comparison of the Galactic longitude profiles of DGE by assuming different two-zone diffusion model parameters. (a) In the energy range of $10-63$ TeV. (b) In the energy range of $63-1000$ TeV. The lines represent the total contribution from the CR-ISM interaction, the LHAASO PWNe/Halo, and the Cygnus bubble. The units of $D_{1}$ are cm2 s-1.